OpenStudy (anonymous):

What is the sum of the geometric sequence –3, 18, –108, … if there are 8 terms? PLEASE HELP!!

OpenStudy (schrodinger):

I'm assuming these are the first three terms in the sequence?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

S_8 = a_1(1-r^n)/1-r so s_8 = -3(1- 6^8)/1-8 3/7(1-6^8) let the scical do the rest :)

OpenStudy (schrodinger):

Scical?

OpenStudy (lgbasallote):

scientific calculator

OpenStudy (schrodinger):

mcast, to clear things up a little bit, the general formula for the sum of a geometric sequence is this: \[a(1-r ^{n})\div(1-r).\]

OpenStudy (anonymous):

it came out to be -1679615.571 when the answer choices are 719,835; -719,835; -119,973; 119,973

OpenStudy (anonymous):

when i did the problem the answer i got was 839808 and that is still not in the choices

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

you there

OpenStudy (schrodinger):

"What is the sum of the geometric sequence –3, 18, –108, … if there are 8 terms? PLEASE HELP!!" Let me see what I can do with this...(By the way, apologies to Igbasallote if i'm doing what you already did) Let's assume this premise: \[s = a(1-r ^{n})\div(1-r)\]. Where a is the first term, s is the sum of the sequence, r is the ratio between terms, and n is the number of terms. Our first term has a value of -3. Therefore, \[a = 3.\] The ratio between terms, \[r\], is equal to the value of the second term divided by the first. or the third divided by the second, or the fourth divided by the third, etc, etc, etc..\[(18)\div(-3) = -6 ... r = -6.\] Since you're aware there are 8 terms, \[n = 8.\] Now, let's plug in to solve:\[s _{8}=(1-(-6^{8}))\div(1-(-6)).\] \[(-6)^{8}=-1679616\] \[(1-(-1679616)) = (1+1679616) = 1679617\] \[(1-(-6)) = (1+6) = 7.\] \[(1679617)\div(7)=239945.286\] Now if i'm correct, S SHOULD...equal that value.

OpenStudy (schrodinger):

Hey...someone else should check my work, the answers he gave don't match up and i'm pretty sure that what I did was right. Something's fishy.